consider the following bivariate utility function


U=25X^2/5Y^3/5


1.find marginal utility of x and y

1
Expert's answer
2022-02-04T07:47:51-0500

Solution:

"U = 25X^{\frac{2}{5} } Y^\frac{3}{5}"

Marginal Utility of X (MUx) = "\frac{\partial U} {\partial X}"= "10X^{-\frac{3}{5} } Y^\frac{3}{5}"

 

Marginal Utility of Y (MUy) = "\frac{\partial U} {\partial Y}" = "15X^{\frac{2}{5} } Y^{-\frac{2}{5} }"